### v1

# Detection of extended X-ray emission surrounding PSR B0656+14 with eROSITA

Abstract
Extended very-high-energy $\gamma$-ray emission from middle-aged pulsars as revealed recently by several groundbased $\gamma$-ray experiments has strong implication on the transport of high-energy particles in the interstellar medium surrounding those pulsars. The $\gamma$-ray emission is widely believed to be produced by high-energy electrons and positrons accelerated by the pulsar wind nebulae when scattering off the interstellar radiation field via the inverse Compton process. Multiwavelength counterparts of the $\gamma$-ray halos are expected to be present, which are, however, not observed yet. In this work, we report for the first time the detection of extended X-ray emission from $\sim 0.2^{\circ}$ radius region of PSR B0656+14 with eROSITA. The spectrum of the emission can be described by a power-law function with an index of $\sim 3.7$. The fluxes decrease with radius faster than the prediction of the particle diffusion and synchrotron radiation in a uniform magnetic field, suggesting the existence of a radial gradient of the magnetic field strength as $\sim r^{-1}$. The magnetic field strength in the X-ray emitting region is constrained to be $4-10~{}\mu$G.
1 Introduction
The transport of Galactic cosmic rays (CRs) in the interstellar medium (ISM) is a fundamental question of astroparticle physics. Due to the scattering with the random magnetic turbulence, charged CRs propagate diffusively in the Milky Way (Berezinskii et al., 1990; Strong et al., 2007). The average diffusion coefficient, usually parameterized as a function of particle’s rigidity and assumed to be spatially uniform, can be inferred from the measurements of secondary to primary ratios of CRs (Aguilar et al., 2016; Alemanno et al., 2022). Possible spatial variations of the diffusion coefficient are expected, due to the changes of properties of the ISM. However, such variations cannot be directly measured from the observations of CRs. It has been shown that large extended $\gamma$-ray halos at very high energies (VHE) (Abeysekara et al., 2017; Aharonian et al., 2021; Linden et al., 2017; Hooper & Linden, 2022; H. E. S. S. Collaboration et al., 2023b, a; Albert et al., 2023), which are expected to be produced by the inverse Compton scattering between high-energy electrons and positrons accelerated by the central pulsar wind nebulae (PWN) and the interstellar radiation field during the propagation process of electrons and positrons, are very powerful sources to probe the transport properties of particles. Particularly, a very slow diffusion is inferred from the morphologies of these pulsar halos, indicating that the propagation is spatially dependent (Fang et al., 2018; Profumo et al., 2018).
Identifying multi-wavelength counterparts of the VHE pulsar halos are particularly important in understanding the emission mechanism of the halos and constraining the magnetic properties of the surrounding ISM. Possible extended GeV $\gamma$-ray emitting halo around Geminga in the Fermi-LAT data has been reported (Di Mauro et al., 2019). However, a controversial result was given in Xi et al. (2019), making this detection inconclusive. Efforts to search for X-ray and radio counterparts have been paid, but no clear detection has been found yet (Liu et al., 2019a; Khokhriakova et al., 2024; Manconi et al., 2024; Wu et al., 2024; Hooper et al., 2024).
In this work, we report the detection of an extended X-ray halo in a region with radius of $0.2^{\circ}$ around the Monogem pulsar (PSR B0656+14, Right Ascension $=104.951117^{\circ}$, Declination $=+14.239175^{\circ}$), with the eROSITA data taken during its Performance Verification (PV) phase (Schwope et al., 2022). The PWN associated with PSR B0656+14 as revealed by Chandra is about $15^{\prime\prime}$ (Bîrzan et al., 2016), which is much smaller than the emission size found in this work. The most natural explanation of the extended X-ray halo is the synchrotron radiation of electrons and positrons which give rise to the VHE $\gamma$-ray emission. The magnetic field strength is thus constrained to be $4-10~{}\mu$G based on the multi-wavelength data.
2 Data Reduction
The eROSITA observations (OBS_ID: 300000) of PSR B0656+14 were performed on October 14th, 2019 in pointing mode [^1] We didn’t include the data in survey mode from the eROSITA Data Release 1(DR1) due to very limited exposure time and quite different instrumental response. during Calibration and Performance Verification (Cal-PV) phase. Due to the operation, telescope module (TM) 1 failed to function properly, while TM 5 and TM7 were affected by optical light leaking, as they do not equip an aluminium optical light filter (Predehl et al., 2021). To ensure optimal sensitivity in the soft X-ray band and extend the low-energy range down to 0.2 keV, we utilized data only from TM2, TM3, TM4, and TM6, i.e., TM8 products. The data were processed using the eROSITA Science Analysis Software System (eSASS) for DR1. Periods of high background activity were excluded by employing the flaregti task, yielding an effective exposure of 99.99 ks.
For spectral extraction, we defined an annular region centered on PSR B0656+14 with radii from 4 to 26 to avoid central source contamination and minimize uncertainties due to edge effects of the instrument’s field of view. Within this region, approximately 530 point sources were identified and excluded using the ermldet tool (Fig. A1; see Appendix A for more details). This annular region was subsequently divided into 12 concentric rings, each with a radius increment of 2, resulting in 12 spectra and their corresponding response files. When generating the response matrices and auxiliary response files, the parameter exttype and psftype of the tool srctool were specifically set to “tophat” and “none” with assuming the the diffuse emission region are very extended. Each spectrum includes an instrumental background spectrum derived from filter wheel closed data. Ultimately, we obtained 12 spectra spanning an energy range of 0.2–2 keV. These spectra are re-binned to ensure sufficient photon counts per channel for better statistical reliability (see Fig. A2 in Appendix A).
3 Spectral fitting and results
The signal we search for may be spatially non-uniform. However, all these 12 annular regions share the same diffuse background emission. Therefore, we jointly fit these 12 spectra using a common set of background parameters, but allowing the flux of the signal component to vary independently across the rings. The spectrum of the signal component is modelled as a power-law, with a constant slope across the 12 annuli. Its foreground absorption should be similar to that of the central point source. Schwope et al. (2022) analyzed the central source using various models, and estimated the HI column density to be about $(1\sim 2.8)\times 10^{20}\,\rm cm^{-2}$. We thus constrain the foreground column density in a similar range.
The diffuse X-ray background is theoretically a composite of several components (e.g., Ponti et al., 2023). The first is the cosmic X-ray background (CXB) predominantly originating from active galactic nuclei and galaxy clusters. Its spectrum is characterized by a power-law distribution with a fixed slope of $\sim 1.46$ and a fixed normalization of $\sim 11.6\,\rm photons$ $\rm cm^{-2}s^{-1}sr^{-1}keV^{-1}$ at 1 keV. Its foreground absorption is modeled based on the HI column density of the Milky Way. The second background is the emission from the local hot bubble (LHB) and the solar wind charge exchange (SWCX). In CCD spectra, these components can be modeled using an APEC plasma model with a temperature of $\sim$0.1 keV, assuming solar metallicity and no foreground absorption (Ponti et al., 2023). The third one is the emission from the Milky Way’s circum-galactic medium (CGM), Galactic halo, and unresolved stellar contributions. The CGM contains a hot, tenuous plasma, with a characteristic temperature of 0.15-0.2 keV and low metallicity. Some high-temperature plasma around 0.4-0.7 keV in the halo or the Galactic corona and unresolved M-dwarf stars contribute soft X-rays in the 0.6-2 keV band. Consequently, an empirical fitting approach is employed, utilizing two absorbed APEC components for the third background emission: one with an initial temperature around 0.2 keV with free metallicity and the other is fixed at 0.7 keV with solar abundance. We fix the temperature and metallicity of the latter one due to the insufficient constraints of the data as Yeung et al. (2024). Both APEC components (CGM and Galactic halo) and CXB share the same fixed absorption HI column density of $1.28\times 10^{21}\,\rm cm^{-2}$ from the HI observation. In this work, we use the tbabs model (Wilms et al., 2000) to estimate the ISM absorption. Finally, the adopted model of the background components is $\tt(tbabs*(powerlaw_{CXB}+apec_{CGM}+apec_{Halo})+apec_{LHB})$.
Spectral fitting was performed using the Interactive Spectral Interpretation System (ISIS; Houck & Denicola, 2000) [^2] Version 1.6.2; https://space.mit.edu/cxc/isis. The fitting shows that a common spectrum power-law excess with photon index of $\sim 4$ on top of the background is found in the data. The fluxes of the power-law excess vary across all annuli, showing a decrease within $4^{\prime}-10^{\prime}$, reaching the level of zero within $10^{\prime}-18^{\prime}$, and rising again at the outermost annuli beyond $18^{\prime}$. The results are presented in Table A1 and Fig. A3. We expect that the rising behavior at large radii may be due to the contribution of weak point sources which are below the detection limit. As shown in Fig. A4, the surface brightness distributions for sources with relatively high fluxes keep flat across the field of view. However, for weak sources, the detection efficiency starts to drop beyond $16^{\prime}$. After correcting the expected contamination from such weak sources, the surface brightness profile of the excess component becomes flat at large radii, as shown in Fig. 1.
To further verify the detected excess component in the inner three annuli, we use the spectrum of the $13^{\prime}-16^{\prime}$ annulus as data-driven background and re-fit the spectra of the inner three annuli. In this procedure, the instrumental background (i.e. filter wheel closed (FWC) data [^3] https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/FWC_dr1) has been subtracted individually for both the signal and background regions. In addition, the BACKSCAL keywords of the three spectra are adjusted to the BACKSCAL columns to account for the energy-dependent differences in their ARFs relative to the background spectrum. The fitting results of model ${\tt tbabs_{diffuse}*powerlaw_{diffuse}}$ for these three annuli are shown in Fig. 2 and summarized in Table 1. The results are consistent with the joint fitting of the 12 spectra within uncertainties, further supporting the detection of an excess power-law component in the $4^{\prime}-10^{\prime}$ region around PSR B0656+14.
The central pulsar is very bright in X rays. Due to the limited point spread function (PSF) of the instrument, the central source may contaminate the emission in outer regions. We estimate that the contribution from the central pulsar to the detected power-law excess due to the PSF can be safely ignored in the $4^{\prime}-6^{\prime}$ region (with possible contamination being smaller than $10\%$; see Appendix A). Therefore this component is not due to the contamination from the central point source. The results in Fig. 1 suggest the presence of an additional diffuse component in the inner three annuli. While such diffuse emission may exist in the outer annuli, its detection is beyond the sensitivity of the current observation.
Figure 1: Fluxes of excess X-rays in 0.2 - 2.0 keV band of the 12 annuli, after correction of the inefficient detection of weak point sources beyond $16^{\prime}$. The background model assumed is the physically motivated three APEC plus CXB model.
Table 1: Fitting results of the non-thermal diffuse emission in three annuli, assuming the $13^{\prime}-16^{\prime}$ data-driven background.
region $N_{\rm H}$ $\Gamma$ Flux0.22.${}_{0.2-2.0~{}\rm keV}$ $\chi^{2}$/dof
(arcmin) ($10^{20}~{}{\rm cm^{-2}}$) ($10^{-12}~{}{\rm erg~{}cm}^{-2}~{}{\rm s}^{-1}~{}{\rm deg}^{-2}$)
$4-6$ $2.08^{+1.36}_{-1.16}$ $3.69^{+0.40}_{-0.36}$ $11.20^{+4.06}_{-2.67}$ 35.93/54
$6-8$ $3.64^{+1.47}_{-1.01}$
$8-10$ $2.05^{+1.01}_{-0.73}$
Notes: Errors are at 90% confidence level. Columns from left to right are: annulus region, HI column density, photon power-law index, flux in 0.2 - 2.0 keV, and fitting $\chi^{2}$ value divided by the number of degree-of-freedom (dof).
Figure 2: X-ray spectra of the non-thermal component for the $4^{\prime}-6^{\prime}$ (black), $6^{\prime}-8^{\prime}$ (red), and $8^{\prime}-10^{\prime}$ (green) annuli. The background model is the $13^{\prime}-16^{\prime}$ data-driven background. The data are binned accordingly to enable that the signal-to-noise ratio (S/N) of each bin at least 3.
Using the data from eRASS:4 (stacked data from the first four consecutive all-sky surveys), Khokhriakova et al. (2024) searched for extended X-ray emission from 5 middle-aged pulsars, and found no significant diffuse emission. The exposure in their analysis is about 500 s for PSR B0656+14, which is perhaps too short to give a sensitive detection. The flux upper limit obtained is about $1.89\times 10^{-12}$ erg s-1 cm-2 deg-2 in the $0.5^{\circ}-1.0^{\circ}$ region and 0.5 - 2.0 keV band, assuming $\Gamma=2$. Their region of interest is beyond the signal region in this work, and the comparison is not direct. Furthermore, the sensitivity of point source detection is not as high as in this analysis, and the unresolved source population may result in a too high background estimate, which affects the detection of possible excess. Using the XMM-Newton observations, an upper limit of $3.3\times 10^{-14}$ erg s-1 cm-2 in the 0.5 - 5.0 keV band within $400^{\prime\prime}$ to $750^{\prime\prime}$ annulus was derived (Wu et al., 2024). In the same sky region and energy band, we obtain a flux of $2.70^{+0.91}_{-0.69}\times 10^{-14}$ erg s-1 cm-2, which is consistent with the XMM-Newton limit.
4 Interpretation
The non-thermal X-ray excess detected by eROSITA shows a radially decline feature, strongly supporting its association with the central pulsar/PWN instead of the background emission associated with the supernova remnant or other sources. The emission also extends far beyond the PWN of PSR B0656+14. These features indicate that it is very likely to be the counterpart of the VHE gamma-ray halo. We thus use a halo model to describe its X-ray to VHE gamma-ray emission.
High-energy electrons and positrons are accelerated by the PWN, which are then injected into the ISM. These particles propagate in the surrounding ISM, and interact with the background radiation field and magnetic field, producing gamma-ray emission via the inverse Compton scattering (ICS) process and X-ray emission via the synchrotron process. The injection spectrum is assumed to be a super-exponentially cutoff power-law function, $Q(E_{e})=Q_{0}E_{e}^{-\alpha}\exp[-(E_{e}/E_{\rm cut})^{\beta}]$, and the injection rate is assumed to be proportional to the spin-down luminosity of the pulsar, $L_{\rm sd}(t)=L_{0}(1+t/\tau)^{-2}$, where $\tau\approx 10^{4}$ yr is a characteristic time scale of the evolution of the pulsar spin-down luminosity. The diffusion coefficient is assumed to be spatially uniform, energy-dependent, as $D(E_{e})=D_{0}(E_{e}/{\rm GeV})^{\delta}$. Here we do not introduce the non-uniform diffusion (Fang et al., 2018) or anisotropic diffusion (Liu et al., 2019b) since we focus on relatively small spatial scales (within $10^{\circ}$ around the pulsar). The diffusion coefficient is adopted to be $D(100~{}{\rm TeV})=4.5\times 10^{27}$ cm2 s-1 as inferred from VHE gamma-ray observations (Abeysekara et al., 2017), and $\delta=1/3$ is assumed. Electrons and positrons experience strong energy losses during the propagation. We include the synchrotron cooling and the ICS cooling with three gray-body components of the background radiation field (Abeysekara et al., 2017): the cosmic microwave background with temperature $T=2.73$ K and energy density $u=0.26$ eV cm-3, an infrared background with $T=20$ K and $u=0.3$ eV cm-3, and an optical background with $T=5000$ K and $u=0.3$ eV cm-3. The ICS cooling rate is calculated following the new formalism of Fang et al. (2021).
Figure 3: Model predicted X-ray spectrum (left) and surface brightness profile (right) of the non-thermal component, compared with the eROSITA data. Dashed lines are for the constant magnetic field assumption, while solid lines correspond to the assumption of $B\propto r^{-1}$ in the inner pc region around the pulsar.
The injection spectral parameters are tuned to describe both the HAWC gamma-ray data and the eROSITA X-ray data. We find that for $\alpha=1.5$, $E_{\rm cut}=35$ TeV, and a very sharp cutoff with $\beta=10$, the measured results can be reproduced, as shown in Fig. 3 and Fig. A5. The conversion efficiency from pulsar spin-down power to the electron/positron power is about 20%. However, we find that the spectral shapes measured by HAWC and eROSITA are hard to be explained well simultaneously. The eROSITA spectrum is very soft, with $\Gamma\approx 3.7$. To account for such a soft spectrum, we require a sharp cutoff of the injection spectrum. Such an injection spectrum gives also a very soft gamma-ray spectrum (Fig. A5), which seems to be too soft compared with that measured by HAWC (Abeysekara et al., 2017). Improved spectral measurements in both gamma and X rays are crucial to further understanding this issue.
The model predicts a relatively flat radial profile, if a constant magnetic field strength ($B=4.2~{}\mu$G) is assumed, as shown by the dashed in the right panel of Fig. 3. The eROSITA data shows a fast decrease trend of the emission, different from the model prediction. It may indicate that there is a radial gradient of the magnetic field. Assuming that [^4] Here we assume that the magnetic field approaches to a constant ISM value of 2.5 $\mu$G at large distance from the pulsar. $B(r)={\rm max}[3.3(r/{\rm pc})^{-1},2.5]~{}\mu$G, the measured radial profile can be reproduced, as shown by the solid line in Fig. 3.
5 Conclusion and discussion
Extended VHE gamma-ray halos around middle-aged pulsars have attracted great attention in recent years, due to their broad connection with the acceleration of energetic particles in the PWN, the propagation of these particles in the surroundings of pulsars, as well as the properties of the ISM. The multi-wavelength counterparts of the gamma-ray halos have been expected to be present, which are, however, not detected yet in any bands other than VHE gamma rays. In this work, we report the first detection of such a counterpart associated with PSR B0656+14 in the X-ray band, with 100 ks observations carried out by eROSITA. An excess component with a power-law spectrum (spectral index $\Gamma=3.69$) extending up to $10^{\prime}$ away from the central pulsar has been identified in the data. The emission size of the excess is much more extended than the X-ray PWN of PSR B0656+14, indicating that it may be produced by particle escaping from the PWN. The fluxes of the excess emission also show fast decline with the increase of distance from the pulsar, which supports its association with the central object rather than unknown background. The pulsar halo model with a high-energy electron and positron cloud surrounding the pulsar/PWN can account for the X-ray and gamma-ray observations simultaneously. The flux of the X-ray emission and its radial profile give a direct measure of the magnetic field in the vicinity of PSR B0656+14, which we find favors the existence of a radial gradient as $B(r)\propto r^{-1}$.
The spectrum of the X-ray excess is somehow soft compared with the HAWC gamma-ray observations. The current simple isotropic diffusion model seems difficult to explain both the X-ray and gamma-ray spectra well. Improved measurements of the gamma-ray spectrum with higher sensitivity and better energy resolution by e.g., LHAASO (Ma et al., 2022) can better clarify the spectral behavior at VHE band. From the modelling point of view, as shown in Wu et al. (2024), the anisotropic diffusion model may give softer X-ray spectrum than the isotropic model while keeping similar TeV gamma-ray spectra.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Nos. 12220101003, 12321003, 12273112), the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061).
Appendix A Point source contamination
The central point source, PSR B0656+14, is significantly brighter than the non-thermal component we aim to detect. Therefore, it is crucial to accounting for any residual effects from the central source due to the PSF. The eROSITA PSF is complex, clearly depending on both energy and off-axis angle. Fortunately, the PSF has been thoroughly characterized and accurately modeled using shapelet functions (Brunner et al., 2022), with full sets of parameters stored in the calibration files.
Figure A1: Count map in 0.5 - 2.0 keV band of the one-degree diameter region around PSR B0656+14 observed by eROSITA with TM2-7. White ellipses mask identified point sources out. The solid green circles show the boundaries of the annuli divided for spectral analysis, from 4 arcmin to 28 arcmin with an increment of 2 arcmin. PSR B0656+14 is shown for illustration, but also excluded in spectroscopic analysis.
PSR B0656+14 is slightly offset from the optical axis. To account for this, we adopt a position corresponding to an off-axis of $-5^{\prime}$ in both the $x$ and $y$ directions as a reference. The PSF images for this position from TM2, 3, 4 and 6 are combined to examine the fractions of flux contribution at various radii. At energies of 0.3, 1.5, and 3.0 keV, the flux contributions from the point source in the $3^{\prime}-4^{\prime}$ annulus are approximately 0.3%, 0.6%, and 0.8%, respectively. Beyond 4, the flux contribution become negligible according to the PSF model, with possible fluctuations $\lesssim 0.1$%. In the $4^{\prime}-6^{\prime}$ annulus, the flux of the non-thermal component is estimated to be $\sim$1.14of the central point source in 0.2-2.0 keV. Therefore, starting the innermost annulus at a radius of 4 is justified, ensuring that the fitting process is not contaminated by the central source.
To exclude the contributions of other point sources, we utilize the ermldet tool to mask them according to their PSF profiles. In the pipeline, the semi-minor axis of the source masks is typically around $\sim 40^{\prime\prime}$. From the above PSF images, within a radius of 1, the enclosed flux fractions are approximately 96.5%, 94.4%, and 94.3% at 0.3, 1.5, and 3.0 keV, respectively. According to the eROSITA point source catalog for this region, there are six sources whose fluxes are about 1/50 of that of PSR B0656+14. For these sources, we extended the mask radius to 1.2 to ensure that any PSF residuals would not affect the measurement of the non-thermal component. The final count map after source mask is shown in Fig. A1.
Figure A2: X-ray spectrum of each annulus. The red points are the eROSITA observational data with the FWC data subtracted. The cyan, purple, and green lines represent the instrumental background, the sum of X-ray sky background, and the non-thermal excess emission. The black lines are the total model fitting results. Note that here the background model assumed is the physically motivated three APEC plus CXB model.
Table A1: Fitting results of the excess power-law component in each annulus, assuming the physically motivated three APEC plus CXB background model. Fluxes are not corrected for the weak source contamination.
region $\Gamma$ Flux0.22.${}_{0.2-2.0~{}\rm keV}$ $\chi^{2}$/dof
(arcmin) ($10^{-12}~{}{\rm erg~{}cm}^{-2}~{}{\rm s}^{-1}~{}{\rm deg}^{-2}$)
4 - 6   $3.70^{+0.16}_{-0.15}$ $13.28^{+1.31}_{-1.44}$ 2431.286/1611
6 - 8 $3.78^{+0.84}_{-0.93}$
8 - 10 $1.81^{+0.69}_{-0.74}$
10 - 12 $0.00^{+0.49}_{-0.00}$
12 - 14 $0.00^{+0.24}_{-0.00}$
14 - 16 $0.41^{+0.57}_{-0.41}$
16 - 18 $0.07^{+0.42}_{-0.07}$
18 - 20 $0.31^{+0.31}_{-0.28}$
20 - 22 $0.38^{+0.39}_{-0.33}$
22 - 24 $0.73^{+0.35}_{-0.33}$
24 - 26 $1.23^{+0.41}_{-0.42}$
26 - 28 $1.06^{+0.48}_{-0.43}$
Note: Errors are at 90% confidence level. Columns from left to right are: annulus region, photon power-law index, flux in 0.2 - 2.0 keV, and fitting $\chi^{2}$ value divided by the number of degree-of-freedom.
Figure A3: Fluxes of excess X-rays in 0.2 - 2.0 keV band of the 12 annuli without weak source contamination correction, assuming the physically motivated three APEC plus CXB background model.
Figure A4: Summed fluxes in the 0.2–2.0 keV range for three flux bins of point sources detected per unit area at different angular distances from the pulsar. The red dashed line gives the constant fitting to the low-flux results within 16, and the red dotted line is the exponential fitting to the results beyond 16.
Appendix B Weak source correction
Fig. A4 show the surface brightness profiles of detected sources, for three flux bins. We convert the count rate to the flux of these point sources with the energy conversion factor (ECF) and the power-law model used by Brunner et al. (2022). One can see that for relatively bright sources, the detection efficiency is almost uniform in the field of view. However, for the weakest sources, the detection efficiency drops significantly beyond 16. The inefficiency of weak source detection may be the reason of the rising behavior at large radii as shown in Fig. A3. We adopt a correction method to correct such an effect. We fit the surface brightness profile from 4 to 16 using a constant, and the surface brightness profile beyond 16 using an exponential function. The difference between the constant surface brightness and the exponential function is taken as the correction values for fluxes measured beyond 16. The corrected flux distribution is shown in Fig. 1 of the main text, which gives nearly flat distribution at large radii.
Appendix C Multi-wavelength modelling
Using the pulsar halo model described in the main text, we calculate the ICS emission of the same population of electrons and positrons, and compare it with the HAWC data. The left panel of Fig. A5 shows the spectrum within $10^{\circ}$ radius of the pulsar, and the right panel shows the surface brightness profile between 8 and 40 TeV.
Figure A5: Model predicted $\gamma$-ray spectrum (left) and surface brightness profile (right) of the non-thermal component, compared with the HAWC data (Abeysekara et al., 2017).

### v2

@@
-Detection of extended X-ray emission surrounding PSR B0656+14 with eROSITA
+Possible evidence for extended X-ray emission surrounding PSR B0656+14 with eROSITA
 Abstract
-Extended very-high-energy $\gamma$-ray emission from middle-aged pulsars as revealed recently by several groundbased $\gamma$-ray experiments has strong implication on the transport of high-energy particles in the interstellar medium surrounding those pulsars.
-The $\gamma$-ray emission is widely believed to be produced by high-energy electrons and positrons accelerated by the pulsar wind nebulae when scattering off the interstellar radiation field via the inverse Compton process.
-Multiwavelength counterparts of the $\gamma$-ray halos are expected to be present, which are, however, not observed yet.
-In this work, we report for the first time the detection of extended X-ray emission from $\sim 0.2^{\circ}$ radius region of PSR B0656+14 with eROSITA.
-The spectrum of the emission can be described by a power-law function with an index of $\sim 3.7$.
-The fluxes decrease with radius faster than the prediction of the particle diffusion and synchrotron radiation in a uniform magnetic field, suggesting the existence of a radial gradient of the magnetic field strength as $\sim r^{-1}$.
-The magnetic field strength in the X-ray emitting region is constrained to be $4-10~{}\mu$G.
-1 Introduction
+[…]
+Cosmic rays — Gamma-ray observatories — High energy astrophysics — Non-thermal radiation sources — Pulsars
+I Introduction
 The transport of Galactic cosmic rays (CRs) in the interstellar medium (ISM) is a fundamental question of astroparticle physics.
@@
 Particularly, a very slow diffusion is inferred from the morphologies of these pulsar halos, indicating that the propagation is spatially dependent (Fang et al., 2018; Profumo et al., 2018).
-Identifying multi-wavelength counterparts of the VHE pulsar halos are particularly important in understanding the emission mechanism of the halos and constraining the magnetic properties of the surrounding ISM.
-Possible extended GeV $\gamma$-ray emitting halo around Geminga in the Fermi-LAT data has been reported (Di Mauro et al., 2019).
-However, a controversial result was given in Xi et al.
-(2019), making this detection inconclusive.
-Efforts to search for X-ray and radio counterparts have been paid, but no clear detection has been found yet (Liu et al., 2019a; Khokhriakova et al., 2024; Manconi et al., 2024; Wu et al., 2024; Hooper et al., 2024).
-In this work, we report the detection of an extended X-ray halo in a region with radius of $0.2^{\circ}$ around the Monogem pulsar (PSR B0656+14, Right Ascension $=104.951117^{\circ}$, Declination $=+14.239175^{\circ}$), with the eROSITA data taken during its Performance Verification (PV) phase (Schwope et al., 2022).
-The PWN associated with PSR B0656+14 as revealed by Chandra is about $15^{\prime\prime}$ (Bîrzan et al., 2016), which is much smaller than the emission size found in this work.
-The most natural explanation of the extended X-ray halo is the synchrotron radiation of electrons and positrons which give rise to the VHE $\gamma$-ray emission.
-The magnetic field strength is thus constrained to be $4-10~{}\mu$G based on the multi-wavelength data.
-2 Data Reduction
-The eROSITA observations (OBS_ID: 300000) of PSR B0656+14 were performed on October 14th, 2019 in pointing mode [^1] We didn’t include the data in survey mode from the eROSITA Data Release 1(DR1) due to very limited exposure time and quite different instrumental response.
+Identifying multi-wavelength counterparts of the VHE pulsar halos is particularly important in understanding the emission mechanism of the halos and constraining the magnetic properties of the surrounding ISM.
+Possible extended GeV $\gamma$-ray emitting halo around Geminga in the Fermi-LAT data has been reported (Di Mauro et al., 2019), which is, however, subject to debate (Xi et al., 2019).
+Efforts to search for X-ray and radio counterparts have been paid, but no clear detection has been found yet (Liu et al., 2019a; Khokhriakova et al., 2024; Manconi et al., 2024; Wu et al., 2024; Hooper et al., 2024)  [^1] The extended $\gamma$-ray source HESS J1809-193 was recently suggested to be a candidate pulsar halo (H.
+E.
+S.
+S.
+Collaboration et al., 2023b) with an X-ray halo (Li et al., 2023).
+However, this region is very complicated, and the origin of the $\gamma$-ray emission is ambiguous..
+We note that the spectral similarity between the central pulsar and the diffuse emission may result from model degeneracies, since weak components can often be fitted acceptably by multiple models.
+The leakage fraction beyond $4^{\prime}$ is expected to vary with photon energy and should not reproduce the pulsar spectrum.
+Moreover, the normalization values, serving as a proxy for the source flux, differ significantly between the pulsar model and the diffuse emission(Khokhriakova et al., 2025), further suggesting that the diffuse component is not dominated by PSF leakage.
+[…]
+II Data Reduction
+The eROSITA observations (OBS_ID: 300000) of PSR B0656+14 were performed on October 14th, 2019 in pointing mode [^2] We did not include the data in survey mode from the eROSITA Data Release 1 (DR1) due to very limited exposure time and quite different instrumental response.
 during Calibration and Performance Verification (Cal-PV) phase.
 Due to the operation, telescope module (TM) 1 failed to function properly, while TM 5 and TM7 were affected by optical light leaking, as they do not equip an aluminium optical light filter (Predehl et al., 2021).
-To ensure optimal sensitivity in the soft X-ray band and extend the low-energy range down to 0.2 keV, we utilized data only from TM2, TM3, TM4, and TM6, i.e., TM8 products.
+To ensure optimal sensitivity in the soft X-ray band and extend the low-energy range down to 0.2 keV, we utilized data only from TM2, TM3, TM4, and TM6.
 The data were processed using the eROSITA Science Analysis Software System (eSASS) for DR1.
@@
 When generating the response matrices and auxiliary response files, the parameter exttype and psftype of the tool srctool were specifically set to “tophat” and “none” with assuming the the diffuse emission region are very extended.
-Each spectrum includes an instrumental background spectrum derived from filter wheel closed data.
+Each spectrum includes an instrumental background spectrum derived from filter wheel closed (FWC) data.
 Ultimately, we obtained 12 spectra spanning an energy range of 0.2–2 keV.
@@
 A2 in Appendix A).
-3 Spectral fitting and results
+III Spectral fitting and results
 The signal we search for may be spatially non-uniform.
@@
 The first is the cosmic X-ray background (CXB) predominantly originating from active galactic nuclei and galaxy clusters.
-Its spectrum is characterized by a power-law distribution with a fixed slope of $\sim 1.46$ and a fixed normalization of $\sim 11.6\,\rm photons$ $\rm cm^{-2}s^{-1}sr^{-1}keV^{-1}$ at 1 keV.
+Its spectrum is characterized by a power-law distribution with a fixed slope of $\sim 1.46$ and a fixed normalization of $\sim 11.6\,\rm photons$ $\rm cm^{-2}~s^{-1}~sr^{-1}~keV^{-1}$ at 1 keV.
 Its foreground absorption is modeled based on the HI column density of the Milky Way.
@@
 The results are presented in Table A1 and Fig.
-A3.
+A6.
 We expect that the rising behavior at large radii may be due to the contribution of weak point sources which are below the detection limit.
 As shown in Fig.
-A4, the surface brightness distributions for sources with relatively high fluxes keep flat across the field of view.
+A7, the surface brightness distributions for sources with relatively high fluxes keep flat across the field of view.
 However, for weak sources, the detection efficiency starts to drop beyond $16^{\prime}$.
@@
 To further verify the detected excess component in the inner three annuli, we use the spectrum of the $13^{\prime}-16^{\prime}$ annulus as data-driven background and re-fit the spectra of the inner three annuli.
-In this procedure, the instrumental background (i.e.
-filter wheel closed (FWC) data [^3] https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/FWC_dr1) has been subtracted individually for both the signal and background regions.
+In this procedure, the instrumental background, i.e.
+FWC data [^4] https://erosita.mpe.mpg.de/dr1/AllSkySurveyData_dr1/FWC_dr1/, has been subtracted individually for both the signal and background regions.
 In addition, the BACKSCAL keywords of the three spectra are adjusted to the BACKSCAL columns to account for the energy-dependent differences in their ARFs relative to the background spectrum.
@@
 The results are consistent with the joint fitting of the 12 spectra within uncertainties, further supporting the detection of an excess power-law component in the $4^{\prime}-10^{\prime}$ region around PSR B0656+14.
-The central pulsar is very bright in X rays.
+The central pulsar is very bright in X-rays.
 Due to the limited point spread function (PSF) of the instrument, the central source may contaminate the emission in outer regions.
-We estimate that the contribution from the central pulsar to the detected power-law excess due to the PSF can be safely ignored in the $4^{\prime}-6^{\prime}$ region (with possible contamination being smaller than $10\%$; see Appendix A).
-Therefore this component is not due to the contamination from the central point source.
-The results in Fig.
-1 suggest the presence of an additional diffuse component in the inner three annuli.
-While such diffuse emission may exist in the outer annuli, its detection is beyond the sensitivity of the current observation.
+The PSF model based on groundbased calibration (Brunner et al., 2022) is limited in accuracy at large radii.
+We therefore choose a pointing-mode observation of a bright variable star, V* V702 CrA, to test the effect of the PSF leakage.
+Note that, V* V702 CrA was observed at an off-axis angle of $4.5^{\prime}$, which is even bigger than the off-axis angle of PSR B0656+14 ($2.5^{\prime}$).
+Instrumentally, a larger off-axis angle is expected to give an even broader PSF.
+As shown in Fig.
+A4 of Appendix B, the radial profile of V* V702 CrA is more concentrated than that of PSR B0656+14 in the 0.2–2.0 keV band, supporting the presence of the diffuse emission.
+The fractional energy beyond $4^{\prime}$ of V* V702 CrA agrees well with the PSF model prediction.
+We estimate that the contribution from the central pulsar to the detected power-law excess can be neglected in the $4^{\prime}-6^{\prime}$ region, with possible contamination being smaller than 10% given the excess flux is about 1.4% of the central pulsar’s flux (see Appendix A).
+As a comparison, an estimated fraction of $\sim 2.7\%$ of the central source is expected beyond $4^{\prime}$, according to the eROSITA survey-mode PSF (Churazov et al., 2023).
+This indicates that the PSF in the survey mode is noticeably degraded due to the instrumental blurring compared to that in the pointing mode around the optical axis.
+These results indicate that the observed extended excess around PSR B0656+14 is not likely to be an artifact of the PSF.
+While a well-calibrated PSF remains crucial for such analyses, future deeper observations and improved characterization of the eROSITA pointing PSF will help confirm this conclusion.
 Figure 1: Fluxes of excess X-rays in 0.2 - 2.0 keV band of the 12 annuli, after correction of the inefficient detection of weak point sources beyond $16^{\prime}$.
@@
 Table 1: Fitting results of the non-thermal diffuse emission in three annuli, assuming the $13^{\prime}-16^{\prime}$ data-driven background.
-region $N_{\rm H}$ $\Gamma$ Flux0.22.${}_{0.2-2.0~{}\rm keV}$ $\chi^{2}$/dof
+[…]
 (arcmin) ($10^{20}~{\rm cm^{-2}}$) ($10^{-12}~{\rm erg~cm}^{-2}~{\rm s}^{-1}~{\rm deg}^{-2}$)
-$4-6$ $2.08^{+1.36}_{-1.16}$ $3.69^{+0.40}_{-0.36}$ $11.20^{+4.06}_{-2.67}$ 35.93/54
-$6-8$ $3.64^{+1.47}_{-1.01}$
-$8-10$ $2.05^{+1.01}_{-0.73}$
-Notes: Errors are at 90% confidence level.
-Columns from left to right are: annulus region, HI column density, photon power-law index, flux in 0.2 - 2.0 keV, and fitting $\chi^{2}$ value divided by the number of degree-of-freedom (dof).
+464-6 $2.08^{+1.36}_{-1.16}$ $3.69^{+0.40}_{-0.36}$ $11.20^{+4.06}_{-2.67}$ 35.93/54
+686-8 $3.64^{+1.47}_{-1.01}$
+8108-10 $2.05^{+1.01}_{-0.73}$
 Figure 2: X-ray spectra of the non-thermal component for the $4^{\prime}-6^{\prime}$ (black), $6^{\prime}-8^{\prime}$ (red), and $8^{\prime}-10^{\prime}$ (green) annuli.
@@
 The exposure in their analysis is about 500 s for PSR B0656+14, which is perhaps too short to give a sensitive detection.
-The flux upper limit obtained is about $1.89\times 10^{-12}$ erg s-1 cm-2 deg-2 in the $0.5^{\circ}-1.0^{\circ}$ region and 0.5 - 2.0 keV band, assuming $\Gamma=2$.
+The flux upper limit obtained is about $1.89\times 10^{-12}$ erg s-1 cm-2 deg-2 in the $0.5\arcdeg-1.0\arcdeg$ region and 0.5 - 2.0 keV band, assuming $\Gamma=2$.
 Their region of interest is beyond the signal region in this work, and the comparison is not direct.
@@
 In the same sky region and energy band, we obtain a flux of $2.70^{+0.91}_{-0.69}\times 10^{-14}$ erg s-1 cm-2, which is consistent with the XMM-Newton limit.
-4 Interpretation
+These results confirm the robustness of the detected diffuse component and motivate further modeling of its physical origin, as discussed below.
+IV Interpretation
 The non-thermal X-ray excess detected by eROSITA shows a radially decline feature, strongly supporting its association with the central pulsar/PWN instead of the background emission associated with the supernova remnant or other sources.
@@
 The diffusion coefficient is assumed to be spatially uniform, energy-dependent, as $D(E_{e})=D_{0}(E_{e}/{\rm GeV})^{\delta}$.
-Here we do not introduce the non-uniform diffusion (Fang et al., 2018) or anisotropic diffusion (Liu et al., 2019b) since we focus on relatively small spatial scales (within $10^{\circ}$ around the pulsar).
+Here we do not introduce the non-uniform diffusion (Fang et al., 2018) or anisotropic diffusion (Liu et al., 2019b) since we focus on relatively small spatial scales (within 10° around the pulsar).
 The diffusion coefficient is adopted to be $D(100~{\rm TeV})=4.5\times 10^{27}$ cm2 s-1 as inferred from VHE gamma-ray observations (Abeysekara et al., 2017), and $\delta=1/3$ is assumed.
@@
 3 and Fig.
-A5.
+A8.
 The conversion efficiency from pulsar spin-down power to the electron/positron power is about 20%.
@@
 Such an injection spectrum gives also a very soft gamma-ray spectrum (Fig.
-A5), which seems to be too soft compared with that measured by HAWC (Abeysekara et al., 2017).
+A8), which seems to be too soft compared with that measured by HAWC (Abeysekara et al., 2017).
 Improved spectral measurements in both gamma and X rays are crucial to further understanding this issue.
-The model predicts a relatively flat radial profile, if a constant magnetic field strength ($B=4.2~{}\mu$G) is assumed, as shown by the dashed in the right panel of Fig.
-3.
-The eROSITA data shows a fast decrease trend of the emission, different from the model prediction.
-It may indicate that there is a radial gradient of the magnetic field.
-Assuming that [^4] Here we assume that the magnetic field approaches to a constant ISM value of 2.5 $\mu$G at large distance from the pulsar.
-$B(r)={\rm max}[3.3(r/{\rm pc})^{-1},2.5]~{}\mu$G, the measured radial profile can be reproduced, as shown by the solid line in Fig.
-3.
-5 Conclusion and discussion
+[…]
+V Conclusion and discussion
 Extended VHE gamma-ray halos around middle-aged pulsars have attracted great attention in recent years, due to their broad connection with the acceleration of energetic particles in the PWN, the propagation of these particles in the surroundings of pulsars, as well as the properties of the ISM.
@@
 The flux of the X-ray emission and its radial profile give a direct measure of the magnetic field in the vicinity of PSR B0656+14, which we find favors the existence of a radial gradient as $B(r)\propto r^{-1}$.
-The spectrum of the X-ray excess is somehow soft compared with the HAWC gamma-ray observations.
-The current simple isotropic diffusion model seems difficult to explain both the X-ray and gamma-ray spectra well.
-Improved measurements of the gamma-ray spectrum with higher sensitivity and better energy resolution by e.g., LHAASO (Ma et al., 2022) can better clarify the spectral behavior at VHE band.
-From the modelling point of view, as shown in Wu et al.
-(2024), the anisotropic diffusion model may give softer X-ray spectrum than the isotropic model while keeping similar TeV gamma-ray spectra.
+The identification of weak extended emission is subject to uncertainties of the PSF model due to the lack of on-orbit data of the pointing observation mode.
+A better understanding of the on-orbit PSF is very important in confirming the results reported in this work.
+Physically the extended X-ray emission associated with ultra-high energy $\gamma$-ray sources is not unexpected.
+For instance, Suzuki et al.
+(2025) have recently reported the detection of X-ray diffuse emission around the PeVatron microquasar V4641 Sgr with XRISM observation, supporting the possibility of extended high-energy halos associated with different types of compact sources.
+Multi-wavelength observations are important in pinning down the radiation nature of these source.
+The spectrum of the X-ray excess appears softer compared with the HAWC $\gamma$-ray observations.
+The simple isotropic diffusion model thus has difficulty in simultaneously explaining both the X-ray and $\gamma$-ray spectra.
+Improved measurements of the $\gamma$-ray spectrum with higher sensitivity and better energy resolution, e.g., by LHAASO (Ma et al., 2022), will help clarify the spectral behavior at the VHE band.
+From the modelling perspective, as shown in Wu et al.
+(2024), the anisotropic diffusion scenario may produce a softer X-ray spectrum than the isotropic case while keeping similar TeV $\gamma$-ray spectra.
+Future high-resolution X-ray missions such as Athena will also provide crucial constraints on the diffuse emission.
 Acknowledgements
 This work is supported by the National Natural Science Foundation of China (Nos.
-12220101003, 12321003, 12273112), the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No.
+12220101003, 12321003, 12273112) and the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No.
 YSBR-061).
@@
 PSR B0656+14 is shown for illustration, but also excluded in spectroscopic analysis.
-PSR B0656+14 is slightly offset from the optical axis.
-To account for this, we adopt a position corresponding to an off-axis of $-5^{\prime}$ in both the $x$ and $y$ directions as a reference.
-The PSF images for this position from TM2, 3, 4 and 6 are combined to examine the fractions of flux contribution at various radii.
-At energies of 0.3, 1.5, and 3.0 keV, the flux contributions from the point source in the $3^{\prime}-4^{\prime}$ annulus are approximately 0.3%, 0.6%, and 0.8%, respectively.
-Beyond 4, the flux contribution become negligible according to the PSF model, with possible fluctuations $\lesssim 0.1$%.
-In the $4^{\prime}-6^{\prime}$ annulus, the flux of the non-thermal component is estimated to be $\sim$1.14of the central point source in 0.2-2.0 keV.
+PSR B0656+14 is slightly offset ($\sim 2.5^{\prime}$) from the optical axis.
+To provide a conservative upper bound on the expected PSF leakage, we adopt a position corresponding to an off-axis of $5^{\prime}$ in the $x$ and $0^{\prime}$ in the $y$ directions as a reference, where the $5^{\prime}$ is the minimum spatial interval for Shapelet PSF parameters.
+The Shapelet PSF images for this position from TM2, 3, 4 and 6 are combined to examine the fractions of flux contribution at various radii.
+At energies of 0.3, 1.5, and 3.0 keV, the flux contributions from the point source in the $3^{\prime}-4^{\prime}$ annulus are approximately 0.3%, 0.6%, and 0.7%, respectively.
+Beyond 4, the flux contribution become negligible according to the PSF model, with possible contribution $\lesssim 0.1$%.
+In the $4^{\prime}-6^{\prime}$ annulus, the flux of the non-thermal component is estimated to be $\sim$1.4% of the central point source in 0.2-2.0 keV.
 Therefore, starting the innermost annulus at a radius of 4 is justified, ensuring that the fitting process is not contaminated by the central source.
@@
 Fluxes are not corrected for the weak source contamination.
-region $\Gamma$ Flux0.22.${}_{0.2-2.0~{}\rm keV}$ $\chi^{2}$/dof
+[…]
 (arcmin) ($10^{-12}~{\rm erg~cm}^{-2}~{\rm s}^{-1}~{\rm deg}^{-2}$)
@@
 26 - 28 $1.06^{+0.48}_{-0.43}$
-Note: Errors are at 90% confidence level.
-Columns from left to right are: annulus region, photon power-law index, flux in 0.2 - 2.0 keV, and fitting $\chi^{2}$ value divided by the number of degree-of-freedom.
+Appendix B Validation of the PSF
+The possible PSF leakage from the bright central source was one of the major concerns.
+To test the physical origin of the extended emission, we measured the 0.2–2.0 keV surface brightness profile of PSR B0656+14 after excluding all detectable point sources, correcting for vignetting, and normalizing to the innermost radius.
+We then compared its profile with multiple reference baselines, including both observations under the pointing mode and the survey mode models.
+Figure A3: The eROSITA image of V* V702 CrA in 0.2–2 keV, corrected for vignetting and FWC.
+Two sectors show the sky regions used to extract the radial surface brightness profile, with a width of $2^{\prime}$.
+Small circles (in cyan and magenta) indicate the regions used to extract the spectra in Fig.
+A5.
+The chosen reference source is the bright star V* V702 CrA (OBSID: 300017), which was observed in the same pointing mode.
+V* V702 CrA is a variable star of spectral type G2 located in the field of dark cloud TGUH2213P1 (Dobashi et al., 2005), a region with low background and sparse point sources.
+V* V702 CrA was observed at an off-axis angle of $\sim 4.5^{\prime}$, which is slightly bigger than the off-axis angle of PSR B0656+14 ($\sim 2.5^{\prime}$).
+Since the eROSITA PSF naturally degrades and broadens with increasing off-axis angle, the reference star’s PSF is expected to be intrinsically broader than that of the pulsar.
+Two reference regions of $90\arcdeg$ sectors (see Fig.
+A3), as well as the $360^{\circ}$ average, are chosen to derive the radial distributions around V* V702 CrA.
+The results are shown in Fig.
+A4.
+We find that the radial profiles for the two $90^{\circ}$ regions are slightly different, with the results of the green sector being more concentrated.
+The $360^{\circ}$ average profile lies in between of the two selected $90^{\circ}$ sectors.
+This may show that there is slight asymmetry of the PSF for the off-axis observations.
+However, the current data are limited in statistics, and the asymmetric effect cannot be clearly characterized.
+In any case, the observed radial profile of PSR B0656+14 is more extended than that of V* V702 CrA, even if the expected profile should be less extended if PSR B0656+14 is indeed a point source.
+This implies that extended emission around the pulsar exists.
+Quantitatively, the fraction of total 0.2–2.0 keV energy in the $4^{\prime}$$6^{\prime}$ annulus for V* V702 CrA is $(0.6\pm 0.2)\%$ as estimated using the east $90^{\circ}$ sector profile (in orange).
+For the other two profiles of V* V702 CrA, such a fraction should be even smaller.
+For PSR B0656+14, the corresponding fraction is $\sim 1.4\%$, significantly higher than the expectation from the V* V702 CrA profile.
+For comparison, we also show in Fig.
+A4 the analytical 0.4–2.3 keV PSF model for the survey mode from Churazov et al.
+(2023) (red line) and the official source stacking PSF of 0.2–2.3 keV (grey line, retrieved from the eROSITA calibration results [^6] https://erosita.mpe.mpg.de/dr1/eROSITA_technical/flight_calibration and valid only within 4’; Merloni et al., 2024).
+We find that the pulsar’s profile is apparently similar to (slightly narrower than) these survey-mode PSF profiles.
+It may partly reflect a coincidence, given the current limitations in the PSF modelling that the source’s intrinsic extent broadens the profile to mimic the instrumental blurring in survey modes.
+Figure A4: Comparison of the surface brightness profiles, normalized at $0.25\arcmin$.
+Blue: PSR B0656+14 (pointing mode); green and orange: V* V702 CrA (pointing mode; the 90° sectors with the same color as labelled in Fig.
+A3); purple: V* V702 CrA (pointing mode; 360° annulus); red: PSF model in the survey mode from Churazov et al.
+(2023); grey: source stacking PSF in survey mode from Merloni et al.
+(2024).
+Spectra were also extracted from the portions of the $4^{\prime}$$8^{\prime}$ and $14^{\prime}$$18^{\prime}$ radial ranges within the east $90^{\circ}$ sector regions around V* V702 CrA, as shown in Fig.
+A5.
+The baseline spectral model consists of the cosmic X-ray background plus local bubble emission with absorption, i.e., $\tt tbabs*(powerlaw+apec)$, giving $\chi^{2}/\mathrm{dof}=369.02/322$.
+By comparing the two spectra, we find no significant difference in the foreground absorption between the two regions, suggesting that it does not affect the surface brightness profile.
+This result is consistent with the HI observation presented by HI4PI Collaboration et al.
+(2016).
+Adding an additional low-metallicity $\tt apec$ component with a fixed temperature $kT=0.83$ keV (corresponding to the emission from the central V* V702 CrA) does not improve the fit significantly (F-test: $F=3.466$, $p=0.0636$).
+Figure A5: Simultaneous fitting of inner and outer regions around V* V702 CrA.
+The cyan and magenta lines correspond to the spectra extracted from the circular regions of the same colors in Fig.
+A3, respectively.
+These results indicate that the leakage of emission from the central pulsar to the region beyond $4^{\prime}$ is very limited, and the excess around PSR B0656+14 is likely not due to the PSF contamination.
+Appendix C Weak source correction
 Figure A6: Fluxes of excess X-rays in 0.2 - 2.0 keV band of the 12 annuli without weak source contamination correction, assuming the physically motivated three APEC plus CXB background model.
@@
 The red dashed line gives the constant fitting to the low-flux results within 16, and the red dotted line is the exponential fitting to the results beyond 16.
-Appendix B Weak source correction
+The original surface brightness profile of the flux excesses is shown in Fig.
+A6, which gives a rising behavior at large radii.
+We expect that this issue should be due to the contamination of weak sources below the detector sensitivity.
 Fig.
-A4 show the surface brightness profiles of detected sources, for three flux bins.
+A7 show the surface brightness profiles of detected sources, for three flux bins.
 We convert the count rate to the flux of these point sources with the energy conversion factor (ECF) and the power-law model used by Brunner et al.
@@
 However, for the weakest sources, the detection efficiency drops significantly beyond 16.
-The inefficiency of weak source detection may be the reason of the rising behavior at large radii as shown in Fig.
-A3.
 We adopt a correction method to correct such an effect.
@@
 1 of the main text, which gives nearly flat distribution at large radii.
-Appendix C Multi-wavelength modelling
+Appendix D Multi-wavelength modelling
 Using the pulsar halo model described in the main text, we calculate the ICS emission of the same population of electrons and positrons, and compare it with the HAWC data.
 The left panel of Fig.
-A5 shows the spectrum within $10^{\circ}$ radius of the pulsar, and the right panel shows the surface brightness profile between 8 and 40 TeV.
+A8 shows the spectrum within 10°radius of the pulsar, and the right panel shows the surface brightness profile between 8 and 40 TeV.
 Figure A8: Model predicted $\gamma$-ray spectrum (left) and surface brightness profile (right) of the non-thermal component, compared with the HAWC data (Abeysekara et al., 2017).